PLEase do problem 2

For your senior design project one of the problems you had to work on was the yield strength of a steel weld joint created by a robot. Your test 15 weld joints and found they had an average yield strength of 270 MPa with a standard deviation of 25MPa. Give a 95% confidence interval for the true yield strength of the weld joint. (by the way Mpa is megapascals measure of force in metric units. 1MPa =145.038 psi)

2. You were also asked to find the average yield strength of an aluminum weld joint made by the robot. Using the standard deviations from the previous problem find the sample size necessary for the 98% confidence interval for the true yield strength of a robotically welded aluminum joint. Let the confidence interval have Width of 20 MPa.
Answer #1

Ans:

confidence interval width=20

Margin of error,E=20/2=10

z=2.326 for 98% CI

Assume sigma=25 (given a=in previous problem)

Sample size required,n=(2.326*25/10)^2=33.81

n=34